# On the Stability of Graph Convolutional Neural Networks: A Probabilistic Perspective

### v1

@@ 5 Empirical Study on Adversarial Edge Perturbation
Methodology. Our goal of adversarial edge perturbation is to maximize the expected embedding perturbations from graph filters or $L$-layer GCNNs
@@ 5.1 Adversarial embedding attack
Results of Graph filters. We evaluate adversarial edge perturbations on two representative filters: the graph adjacency (low-pass) and the graph Laplacian (high-pass), which capture distinct signal propagation behaviors in graph-based data processing. We apply different adversarial edge perturbation methods with a fixed budget of $m=20$ edges (less than 2% of the maximum possible edges in most graphs). The results, shown in Figure 3, are presented as violin plots, illustrating the distribution of embedding perturbations across multiple graph signals—100 randomly sampled signals for synthetic datasets and 18 real-world measurement signals for ENZYMES. This provides a more detailed view of how edge perturbations affect different signal realizations. Across all datasets and filter types, our distribution-aware method Prob-PGD consistently induces the largest overall embedding perturbations, highlighting the effectiveness and robustness of our methodology.
@@
Results of multilayer GCN. We conduct experiments on multilayer GCN architectures [26] using the normalized adjacency matrix $g(\mathbf{S})=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$. The tests are performed on contextual stochastic block models (cSBM) with 100 vertices, where each algorithm perturbs 20 edges. Since the learnable parameters $\{\mathbf{\Theta}^{(l)}\}_{l=1}^{L}$ also influence the final embeddings, we perform 200 repeated experiments with model weights independently resampled from a standard Gaussian distribution to ensure statistical reliability. Figure 4 presents the Frobenius norm of the embedding perturbations at each layer, visualized as box plots over the 200 trials. Across all layers and activation functions, our distribution-aware method Prob-PGD consistently outperforms the baselines, demonstrating the effectiveness of our probabilistic formulation in deeper architectures. We also observe that the performance gap between attack methods narrows in deeper layers, possibly due to the loosening of our theoretical bound with depth and inherent phenomena in deep GCNs, such as oversmoothing. Additional results with different model architectures are provided in Appendix C.4.
@@ 5.2 Adversarial attack on downstream tasks
We further evaluate how embedding-level adversarial perturbations affect the downstream performance of pre-trained GCNN models. Our experiments include three popular architectures: GCN [26], GIN [48], and SGC [47], across different downstream tasks. More details about model architecture are provided in Table 3 in Appendix A.2, and training details are given in Appendix C.2.
Graph classification. We evaluate graph classification performance on two benchmark datasets: MUTAG and ENZYMES. MUTAG contains 188 molecular graphs, divided into two classes based on their mutagenic effect on a bacterium. ENZYMES consists of 600 graphs of protein tertiary structures, categorized into six EC enzyme classes. Graphs from both datasets include numerical node features (e.g., atom types in MUTAG, physical measurements in ENZYMES), and we compute the node feature second-moment matrix from the empirical distribution using $\mathbf{K}_{\text{samp}}=\sum_{i=1}^{d}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$. We pre-train the GCNNs on the training set and keep their parameters fixed during evaluation. To assess prediction robustness, we perturb $5\%$ edges in test set graphs using different attack algorithms and measure classification accuracy before and after perturbation.
Node classification. We study node classification on the Cora dataset, which consists of a single citation graph with 2708 nodes representing scientific publications, categorized into seven classes. Each node is associated with a sparse bag-of-words feature vector, and we compute the second-moment matrix from the empirical distribution. We pre-train GCNNs on the unperturbed graphs with 140 labeled nodes. To assess the robustness of pre-trained models, we apply different algorithms to perturb 1000 edges to the citation graph and evaluate classification accuracy on test set nodes before and after perturbation.
Results. We report classification accuracy and embedding perturbation magnitude (Frobenius norm) in Table 1. Across datasets and architectures, Prob-PGD consistently outperforms both baselines, produces the largest embedding perturbations and greater performance degradation in eight out of nine settings. On MUTAG, for instance, the accuracy drops are particularly significant for GCN and SGC, highlighting the ability of our distribution-aware approach to identify impactful perturbations. We further illustrate in Figure 5 the performance variation of GCN on the MUTAG dataset under varying levels of edge perturbation. The left plot shows that Prob-PGD produces the largest embedding shifts as edge perturbation increases. The right plot demonstrates that embedding-level perturbations effectively transfer to downstream performance degradation. Notably, even small perturbations (1–3%) lead to substantial performance drop, underscoring the importance of robustness analysis.
Dataset Method Prediction Accuracy Embedding Pert. $\|\cdot\|_{F}$
@@
MUTAG Unperturbed 0.79 0.74 0.79 - - -
Random 0.74 0.77 0.81 176 1154 105
Wst-PGD 0.74 0.34 0.74 134 2235 87
Prob-PGD 0.29 0.82 0.34 205 3262 155
ENZYMES Unperturbed 0.56 0.57 0.45 - - -
Random 0.46 0.45 0.42 4845 11273 2458
Wst-PGD 0.50 0.41 0.42 3498 29156 1817
Prob-PGD 0.45 0.37 0.37 5282 39716 3046
Cora Unperturbed 0.81 0.72 0.80 - - -
Random 0.79 0.71 0.78 90 1088 57
Wst-PGD 0.79 0.72 0.78 92 17719 70
Prob-PGD 0.78 0.53 0.77 100 54654 66
@@
Figure 5: Classification performance of a two-layer GCN on the MUTAG dataset under varying levels of edge perturbation.
@@ A.2 Concept definitions
For a general graph filter $g(\mathbf{S})$, the corresponding filter perturbation induced by $\mathcal{P}$ can be computed by first constructing $g(\mathbf{S})$ and $g(\mathbf{S}_{p})$ from the original graph $\mathcal{G}$ and the perturbed graph $\mathcal{G}_{p}$ separately, and then taking their difference $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$. For example, if we consider a polynomial graph filter of the form $g(\mathbf{S})=\sum_{k=1}^{K}\mathbf{A}^{k}$, then the filter perturbation becomes $\mathbf{E}_{g}=\sum_{k=1}^{K}\mathbf{A}^{k}-\mathbf{A}_{p}^{k}$.
@@
Figure 6:
@@
GIN [48] $\tilde{\mathbf{A}}$ $\mathbf{A}-\mathbf{A}_{p}$ max degree
@@ B.5 Proof of Proposition 1
(1) $g(\mathbf{S})=\mathbf{A}$.
@@
(2) $g(\mathbf{S})=\mathbf{L}$
@@ C.2 Experiment setup
Training GCNN models. For graph classification tasks using the ENZYMES and MUTAG datasets, we partition the dataset into training and test sets using an 80%, 20% split. We pre-train graph convolutional neural networks (GCNNs), including two-layer GCN [26], two-layer GIN [48], and two-layer Simple Graph Convolution (SGC) [47], on the original, unperturbed graphs. Each GCNN has a hidden dimension of 64 and uses max-pooling as the final layer. In the adversarial attack experiments, the number of edge perturbations is limited to at most $5\%$ of the total possible edge modifications, i.e., $5\%n^{2}$, where $n$ is the number of nodes in the graph. Each attack is repeated five times on the clean graph, and we report the average classification accuracy.
For node classification, we pre-train the same set of two-layer GCNNs on the original, unperturbed graph using 140 labeled nodes for training and 1000 labeled nodes for testing, following the setup in [26]. The key hyperparameters are set as follows: a hidden dimension of 64, a learning rate of 0.001, and the ReLU function as the activation function. In the adversarial attack experiments, each algorithm perturbs 1000 edges. We repeat the experiments five times and report the average classification accuracy.
All experiments were conducted on an NVIDIA V100 GPU with 16GB of memory. Pretraining of the GCNNs on the input graphs completed within a few minutes. Adversarial attacks on the ENZYMES and MUTAG datasets finished within a few minutes, while attacks on the Cora dataset required slightly more time.
C.3 Comparison with the average-case metric for graph filters
@@
C.4 Embedding perturbation of multilayer GCNNs
@@
C.5 Computing the second-moment matrix $\mathbf{K}$

### v3

@@ 5 Empirical Study on Adversarial Edge Perturbation
Methodology. We consider a task-agonistic edge attack where our goal of adversarial edge perturbation is to maximize the expected embedding perturbations from graph filters or $L$-layer GCNNs
@@ 5.1 Adversarial embedding attack
Results of graph filters embedding. We evaluate adversarial edge perturbations on two representative filters: the graph adjacency (low-pass) and the graph Laplacian (high-pass), which capture distinct signal propagation behaviors in graph-based data processing. We apply different adversarial edge perturbation methods with a fixed budget of $m=20$ edges (less than 2% of the maximum possible edges in most graphs). The results, shown in Figure 3, are presented as violin plots, illustrating the distribution of embedding perturbations across multiple graph signals—100 randomly sampled signals for synthetic datasets and 18 real-world measurement signals for ENZYMES. This provides a more detailed view of how edge perturbations affect different signal realizations. Across all datasets and filter types, our distribution-aware method Prob-PGD consistently induces the largest overall embedding perturbations, highlighting the effectiveness and robustness of our methodology.
@@
Results of multilayer GCN embedding. We conduct experiments on multilayer GCN architectures [26] using the normalized adjacency matrix $g(\mathbf{S})=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$. The tests are performed on contextual stochastic block models (cSBM) with 100 vertices, where each algorithm perturbs 20 edges. Since the learnable parameters $\{\mathbf{\Theta}^{(l)}\}_{l=1}^{L}$ also influence the final embeddings, we perform 200 repeated experiments with model weights independently resampled from a standard Gaussian distribution to ensure statistical reliability. Figure 4 presents the Frobenius norm of the embedding perturbations at each layer, visualized as box plots over the 200 trials. Across all layers and activation functions, our distribution-aware method Prob-PGD consistently outperforms the baselines, demonstrating the effectiveness of our probabilistic formulation in deeper architectures. We also observe that the performance gap between attack methods narrows in deeper layers, possibly due to the loosening of our theoretical bound with depth and inherent phenomena in deep GCNs, such as oversmoothing. Additional results with different model architectures are provided in Appendix C.5.
@@ 5.2 Adversarial attack on downstream tasks
We further evaluate how embedding-level adversarial perturbations affect the downstream performance of pre-trained GCNN models. Our experiments include three popular architectures: GCN [26], SGC [49], and CIN [^2] We modify the standard GIN architecture by replacing the MLP in each layer with a single-layer perceptron, resulting in a simplified convolutional variant of GIN. Experiments (Appendix C.6) show that standard GINs with deeper MLPs exhibit a similar response to edge perturbations as their convolutional counterparts. [50]. More details about model architecture are provided in Table 4 in Appendix A.2, and training details are given in Appendix C.2.
Graph classification. We evaluate classification performance on two benchmark datasets: MUTAG (188 molecular graphs divided into two classes) and ENZYMES (600 protein graphs labeled by six enzyme classes). Both datasets provide numerical node features (e.g., atom types and physical measurements), from which we compute the second-moment matrix as $\mathbf{K}_{\text{samp}}=\sum_{i=1}^{d}\mathbf{x}_{i}\mathbf{x}_{i}^{T}$. For each dataset, we use 80% of the graphs to pre-train 10 GCNNs and keep their parameters fixed during evaluation. To assess their prediction robustness, we perturb $5\%$ edges in test set graphs using different attack algorithms and measure classification accuracy before and after perturbation.
Node classification. We use the Cora citation network (2708 vertices belonging to seven categories) with sparse bag-of-words node features. We pre-train 10 GCNNs on the unperturbed graphs with 140 labeled nodes and keep the model parameters fixed during evaluation. To assess the robustness of these pre-trained models, we apply different algorithms to perturb 1000 edges to the graph and evaluate classification accuracy on test set nodes before and after perturbation.
Dataset Method Prediction Accuracy(%) Embedding Pert. $\|\cdot\|_{F}$
@@
MUTAG Unpert. 70.00±$\pm 2.68$ 82.89±$\pm 4.60$ 68.95±$\pm 1.58$
Random 73.16±$\pm 1.05$ 70.53±$\pm 4.82$ 72.89±$\pm 1.21$ 2.24±$\pm 0.15$ 28.62±$\pm 2.34$ 3.57±$\pm 0.17$
Wst-PGD 69.21±$\pm 2.37$ 76.32±$\pm 8.15$ 68.16±$\pm 1.42$ 1.95±$\pm 0.13$ 37.47±$\pm 9.47$ 3.11±$\pm 0.16$
Prob-PGD 42.89±$\pm 4.86$ 76.58±$\pm 8.02$ 41.58±$\pm 3.87$ 3.09±$\pm 0.16$ 96.26±$\pm 11.39$ 5.21±$\pm 0.25$
ENZYMES Unpert. 56.33±$\pm 3.10$ 56.75±$\pm 2.99$ 54.75±$\pm 3.48$
Random 49.58±$\pm 3.93$ 45.83±$\pm 2.64$ 45.50±$\pm 2.94$ 15.3±$\pm 0.7$ 81.3±$\pm 11.3$ 36.4±$\pm 1.2$
Wst-PGD 52.58±$\pm 3.56$ 39.75±$\pm 3.29$ 48.75±$\pm 4.27$ 11.5±$\pm 0.5$ 264.9±$\pm 34.7$ 27.2±$\pm 0.9$
Prob-PGD 44.25±$\pm 2.09$ 36.75±$\pm 4.62$ 40.75±$\pm 3.83$ 17.9±$\pm 0.9$ 297.9±$\pm 57.1$ 43.6±$\pm 1.6$
Cora Unpert. 80.06±$\pm 0.51$ 74.92±$\pm 1.86$ 80.33±$\pm 0.29$
Random 78.30±$\pm 0.82$ 71.75±$\pm 1.85$ 78.12±$\pm 0.64$ 93±$\pm 4$ 974±$\pm 496$ 136±$\pm 4$
Wst-PGD 78.47±$\pm 0.44$ 73.76±$\pm 2.07$ 78.30±$\pm 0.33$ 97±$\pm 7$ 16901±$\pm 12495$ 151±$\pm 5$
Prob-PGD 77.88±$\pm 0.33$ 56.67±$\pm 4.57$ 77.49±$\pm 0.29$ 106±$\pm 5$ 43565±$\pm 33035$ 160±$\pm 6$
@@
Figure 5: Classification performance of a pre-trained two-layer GCN on the MUTAG dataset under varying levels of edge perturbation.
Results. We report classification accuracy and embedding perturbation magnitude (measured by the Frobenius norm) in Table 1, averaged over 10 pre-trained models. Across datasets and architectures, Prob-PGD consistently induces the largest embedding perturbations, which result in greater performance degradation in eight out of nine cases. Figure 5 further illustrates the performance of a pre-trained two-layer GCN on the MUTAG dataset under varying levels of edge perturbation. The left plot shows that Prob-PGD consistently induces the largest embedding perturbations across all perturbation levels. The right plot demonstrates that embedding-level perturbations transfer to downstream performance degradation. Notably, when the edge perturbation exceeds 3%, Prob-PGD reduces model accuracy to a level comparable to a trivial classifier that always predicts the majority class. Overall, our proposed algorithm prob-PGD, although task-agnostic, still causes significant performance degradation in pre-trained GCNNs across different tasks. These results underscore the vulnerability of GCNNs to even small edge perturbations and reinforce the importance of comprehensive stability analysis in such settings.
@@ A.2 Concept definitions
For a general graph filter $g(\mathbf{S})$, the corresponding filter perturbation induced by $\mathcal{P}$ can be computed by first constructing $g(\mathbf{S})$ and $g(\mathbf{S}_{p})$ from the original graph $\mathcal{G}$ and the perturbed graph $\mathcal{G}_{p}$ separately, and then taking their difference $\mathbf{E}_{g}=g(\mathbf{S})-g(\mathbf{S}_{p})$. We summarize in Table 3 some common filters and their perturbation.
Graph Filter $g(\mathbf{S})$ Filter Perturbation $\mathbf{E}_{g}$
polynomial of adjacency matrix: $\sum_{j=1}^{k}c_{j}\mathbf{A}$ $\sum_{j=1}^{k}c_{j}(\mathbf{A}-\mathbf{A}_{p})$
polynomial of Laplacian: $\sum_{j=1}^{k}c_{j}\mathbf{L}$ $\sum_{j=1}^{k}c_{j}(\mathbf{L}-\mathbf{L}_{p})$
low-pass filter: $(\mathbf{I}+\alpha\mathbf{L})^{-1}$ $(\mathbf{I}+\alpha\mathbf{L})^{-1}-(\mathbf{I}+\alpha\mathbf{L}_{p})^{-1}$
heat diffusion filter: $e^{-\tau\mathbf{L}}$ $e^{-\tau\mathbf{L}}-e^{-\tau\mathbf{L}_{p}}$
Table 3: Common graph filters from [39, 37] and their perturbations
@@
Figure 6: Attributed graph and graph perturbation.
@@
GIN [^3] We modify the standard GIN architecture by replacing the MLP in each layer with a single-layer perceptron, resulting in a simplified convolutional variant of GIN. [50] $(1+\epsilon^{(k)})\mathbf{I}+\mathbf{A}$ $\mathbf{A}-\mathbf{A}_{p}$ max degree
@@ B.5 Proof of Proposition 1
(Case 1.) Proof for adjacency filter $g(\mathbf{S})=\mathbf{A}$.
@@
(Case 2.) Proof for Laplacian filter $g(\mathbf{S})=\mathbf{L}$
@@ C.2 Experiment setup
Training GCNN models. For graph classification tasks using the ENZYMES and MUTAG datasets, we partition the dataset into training and test sets using an 80%, 20% split. We pre-train graph convolutional neural networks (GCNNs), including two-layer GCNs [26], two-layer SGCs [49], and two-layer GIN, where GIN refers to a specialized variant of the original model in which the MLPs are replaced with single-layer perceptrons. All models use a hidden dimension of {32,64}, followed by a ReLU activation, and employ max-pooling as the final readout layer. Training on ENZYMES and MUTAG both use the standard cross-entropy loss. Models are optimized using the Adam optimizer with learning rates between $0.005$ and $0.01$, weight decay $10^{-3}$, and 50 to 100 training epochs. In the adversarial attack experiments, the number of edge perturbations is limited to at most $5\%$ of the total possible edge modifications, i.e., $5\%n^{2}$, where $n$ is the number of nodes in the graph.
For node classification, we pre-train two-layer GCNs [26], two-layer SGCs [49], and two-layer SUM-GCNs on the original, unperturbed graph, using 140 labeled nodes for training and 1,000 nodes for testing, following the standard setup in [26]. Each GCNN has a hidden dimension of 64 and uses ReLU as the activation function. The models are trained using a standard cross-entropy loss over the labeled training nodes. Training was conducted using the Adam optimizer with a learning rate of 0.01 and weight decay $5\times 10^{-4}$ for 50 epochs. In the adversarial attack experiments, each algorithm perturbs 1000 edges. We repeat the experiments five times and report the average classification accuracy.
All experiments were conducted on an NVIDIA V100 GPU with 16GB of memory. Pretraining of the GCNNs on the input graphs was completed within a few minutes. Adversarial attacks on the ENZYMES and MUTAG datasets finished within a few minutes, while attacks on the Cora dataset required slightly more time due to the larger graph size, where Prob-PGD and Wst-PGD both take nearly 10 minutes to complete, with the maximum number of PGD iterations set to 250.
C.3 Computing the second-moment matrix $\mathbf{K}$
C.4 Comparison with the average-case metric for graph filters
@@
C.5 Embedding perturbation of multilayer GCNNs
@@
C.6 Experiments on message passing neural networks
This study mainly focuses on analyzing the stability of Graph Convolutional Neural Networks (GCNNs), which represent one of the two main paradigms in graph neural networks. The other major paradigm consists of message passing neural networks (MPNNs). While both frameworks aggregate node information based on the underlying graph, MPNNs do not necessarily rely on graph filter convolutions. One notable example of MPNN is the Graph Isomorphism Network (GIN) [50]. In GIN, vertex embeddings are computed recursively. For a given vertex $v$, its embedding at layer $l$ is obtained by aggregating information from its neighbors as well as from itself at the previous layer, according to the following:
$\mathbf{x}_{v}^{(l)}=\text{MLP}^{(l)}\left(\left(1+\epsilon^{(l)}\right)\cdot\mathbf{x}_{v}^{(l-1)}+\sum_{u\in\mathcal{N}(v)}\mathbf{x}_{u}^{(l-1)}\right)$
where $\mathcal{N}(v)$ denotes the set of neighbors of vertex $v$ and $\epsilon^{(l)}$ is either a learnable parameters or a fixed scalar. When the multi-layer perceptron (MLP) is reduced to a single-layer perceptron, the model simplifies to a GCNN with a graph filter of the form $\mathbf{A}+\epsilon^{(l)}\mathbf{I}$. Empirical studies on the stability of GIN with a single-layer perceptron are presented in Section 5. Here, we extend this investigation to GIN architectures with a two-layer perceptron (GIN-2) and a three-layer perceptron(GIN-3), both implemented within a two-layer GIN framework. Using the same experimental settings as in Section 5, we pre-train 10 GINs on unperturbed graphs. Table 5 reports the prediction accuracy and embedding perturbation under various edge perturbation algorithms. Across different datasets and models, Prob-PGD induces the largest embedding perturbations, and further results in greater performance degradation for ENZYMES and Cora classification. Overall, GIN models with deeper MLPs exhibit a similar response to edge perturbations as their single-layer counterparts. This observation suggests that our filter-based analysis might be extended to study the stability of MPNNs by investigating graph filters that capture their underlying message passing mechanism.
Dataset Method Prediction Accuracy(%) Embedding Pert. $\|\cdot\|_{F}$
GIN-2 GIN-3 GIN-2 GIN-3
MUTAG Unpert. 82.89$\pm$4.60 82.11$\pm$5.49 - -
Random 70.53$\pm$4.82 70.53$\pm$5.37 28.62$\pm$2.34 28.43$\pm$3.79
Wst-PGD 76.32$\pm$8.15 76.32$\pm$9.04 37.47$\pm$9.47 37.37$\pm$11.16
Prob-PGD 76.58$\pm$8.02 77.11$\pm$7.81 96.26$\pm$11.39 93.47$\pm$10.32
ENZYMES Unpert. 58.92$\pm$3.59 50.08$\pm$5.30 - -
Random 48.25$\pm$3.75 41.83$\pm$4.15 92.44$\pm$12.56 28.35$\pm$5.16
Wst-PGD 39.33$\pm$4.83 41.00$\pm$3.14 295.62$\pm$35.02 74.06$\pm$12.83
Prob-PGD 37.00$\pm$3.25 35.58$\pm$3.27 345.54$\pm$49.35 83.59$\pm$20.24
Cora Unpert. 73.76$\pm$1.95 70.72$\pm$2.30 - -
Random 69.73$\pm$2.63 66.95$\pm$3.16 1129$\pm$344 1498$\pm$1127
Wst-PGD 72.55$\pm$2.25 69.19$\pm$2.20 16529$\pm$7638 11105$\pm$6508
Prob-PGD 57.95$\pm$3.35 54.09$\pm$5.03 40343$\pm$23605 25784$\pm$13428
Table 5: Prediction accuracy of GINs under edge perturbations produced by different methods.